# Noble gravitational atoms: Self-gravitating black hole scalar wigs with angular momentum number

### v1

@@ 2.2 The $\ell$-boson star ansatz
For the matter content, we consider a family of $2\ell+1$ classical complex scalar fields $\Phi_{m}$, each with the same mass $\mu$, which gives rise to the stress-energy tensor
@@
Initial data for $(\phi_{\ell},\pi_{\ell})$ on a $t={\rm const.}$ hypersurface can be freely specified. They determine the relevant components (8) of the stress-energy tensor, up to the mass function $m$ which appears in the factor $\gamma$ in $\chi_{\ell}=\gamma^{-1}\phi_{\ell}^{\prime}$. In turn, $m$ can be obtained by integrating equation (3b). This allows one to define the three-metric and extrinsic curvature by means of equations (4), and by construction they automatically satisfy the constraints. In what follows, we describe a method for obtaining exact initial data approximating giving rise to quasi-stationary solutions of the evolution system (3) and (9).
@@ 4 Conclusions
We have presented new astrophysically relevant solutions to the Einstein-Klein-Gordon equations, describing quasi-stationary systems that contain a black hole surrounded by a self-gravitating scalar field distribution. Although spherically symmetric, these objects are parametrized by an angular momentum number $\ell$, and therefore they have been named “noble gravitational atoms”. Noble gravitational atoms can vary greatly in shape, size, mass, and lifetime, depending on the following parameters: the angular momentum number $\ell$, the scalar field mass $\mu$, the black hole mass $M$, and the scalar field amplitude $A$.
@@
In general, noble gravitational atoms, as their $\ell$-boson star counterparts, can become larger as the angular momentum number $\ell$ increases. This characteristic may help extend the galactic halo core model mentioned earlier to cover a larger fraction of the galaxy, more so if one also considers solutions that combine multiple angular momentum numbers $\ell$ (and possibly even multiple mode numbers $n$ beyond the ground state), potentially allowing the model to reproduce the full galactic size. We leave such studies for future work.

### v2

@@ 2.2 The $\ell$-boson star ansatz
For the matter content, we consider a family of $2\ell+1$ classical complex scalar fields $\Phi_{m}$, each with the same mass $\mu$, which gives rise to the stress-energy tensor [^1] Although some forms of interactions could in principle be included in our model, we only consider free scalar fields in this work. Self-interactions may be relevant for axion-like candidates; however, they are generically controlled by the axion mass and are therefore extremely weak in the ultralight case. Moreover, such interactions involve higher powers of the axion field in the equations of motion. Since the scalar field remains bounded in our configurations, there always exists a region of the solution space for which the self-interaction effects remain negligible. For these reasons we neglect self-interaction terms in our work.
@@
Initial data for $(\phi_{\ell},\pi_{\ell})$ on a $t={\rm const.}$ hypersurface can be freely specified. They determine the relevant components (8) of the stress-energy tensor, up to the mass function $m$ which appears in the factor $\gamma$ in $\chi_{\ell}=\gamma^{-1}\phi_{\ell}^{\prime}$. In turn, $m$ can be obtained by integrating equation (3b). This allows one to define the three-metric and extrinsic curvature by means of equations (4), and by construction they automatically satisfy the constraints. In what follows, we describe a method for obtaining exact initial data giving rise to approximate quasi-stationary solutions of the evolution system (3) and (9).
@@ 4 Conclusions
We have presented new astrophysically relevant approximate solutions to the Einstein-Klein-Gordon equations, describing quasi-stationary systems that contain a black hole surrounded by a self-gravitating scalar field distribution. Although spherically symmetric, these objects are parametrized by an angular momentum number $\ell$, and therefore they have been named “noble gravitational atoms”. Noble gravitational atoms can vary greatly in shape, size, mass, and lifetime, depending on the following parameters: the angular momentum number $\ell$, the scalar field mass $\mu$, the black hole mass $M$, and the scalar field amplitude $A$.
@@
In general, noble gravitational atoms, as their $\ell$-boson star counterparts, can become larger as the angular momentum number $\ell$ increases. This characteristic may help extend the galactic halo core model mentioned earlier to cover a larger fraction of the galaxy, more so if one also considers solutions that combine multiple angular momentum numbers $\ell$ (and possibly even multiple mode numbers $n$ beyond the ground state), potentially allowing the model to reproduce the full galactic size.
In order to asses whether noble gravitational atoms can be regarded as viable astrophysical candidates, there is a series of open questions that remain to be addressed. A key aspect is their stability. $\ell$-boson stars are known to be stable against radial perturbations [43, 44, 49], but they have been shown to be unstable against generic (non-radial) perturbations for $\ell>1$, at least in the nonrelativistic regime [50]. Similar behavior could be expected for noble gravitational atoms; however, a detailed stability analysis is required to establish this conclusively.
Another important aspect of these new configurations concerns their gravitational wave emissions. Noble gravitational atoms are expected to exhibit gravitational spectra with distinctive features that differ from those of simple vacuum black hole solutions. The role of the angular momentum parameter $\ell$ has been explicitly studied in simulations of binary $\ell$-boson star mergers [47], where it was shown to affect the emitted radiation. Such distinctive signatures may be observable with current and future gravitational wave detectors [58], potentially providing a mechanism to prove the existence of $\ell$ scalar field distributions. A detailed investigation of these signals is left for future work.
Finally, we emphasize that the specific value $m\sim 10^{-22}\rm{eV}$ used for the results presented in our figures should be regarded purely as a reference choice. Current constraints on ultralight dark matter are in tension when different cosmological and astrophysical observations are combined (see, e.g., figure 18 in reference [19]). On the one hand, cosmological probes such as the Lyman-alpha forest, the cosmic microwave background, and large-scale structure typically require masses of the order of $m\gtrsim 10^{-22}-10^{-21}\rm{eV}$ if the ultralight field is to constitute the dominant component of dark matter, in order to avoid excessive suppression of small-scale structure [59]. On the other hand, fits to the internal dynamics of dwarf spheroidal galaxies, which are dominated by dark matter, often favor lighter masses, $m\sim 10^{-22}\rm{eV}$, or even below, motivated by the ability of wave-like effects to alleviate the cusp profiles suggested by $N$-body simulations in the context of standard cold dark matter [60].
As a consequence, there is currently no single mass range that simultaneously satisfies all observational constraints within the simplest realizations of ultralight dark matter, and the precise bounds remain highly model dependent. Moreover, many of these constraints rely on specific assumptions regarding the structure and equilibrium profiles of dark matter halos, with most galactic analyses focusing on the standard spherically symmetric $\ell=0$ ground-state configuration. Noble gravitational atoms, by contrast, correspond to distinct equilibrium states, and their density and kinematic profiles can differ substantially from those usually assumed. This opens the possibility that the corresponding phenomenological bounds on the boson mass may be modified, potentially reconciling cosmological and astrophysical constraints. Assessing this in detail would require dedicated fits to observational data using the profiles studied here, which lies beyond the scope of the present work.